# ADPro: a Test-time Adaptive Diffusion Policy for Robot Manipulation via Manifold and Initial Noise Constraints

### v1

@@ Abstract
1 Introduction
Autonomous robotic manipulation in unstructured environments requires policies that not only predict accurate actions but also generalize well across diverse tasks and scenes. Early approaches based on rule-based controllers  (Miller and Allen 2004; Li, Fu, and Pollard 2007; Dang, Weisz, and Allen 2011) or dense discriminative models  (Ten Pas et al. 2017; Liang et al. 2019; Mousavian, Eppner, and Fox 2019; Murali et al. 2020; Fang et al. 2020; Gou et al. 2021; Zhao et al. 2021; Depierre, Dellandréa, and Chen 2021; Wei et al. 2021; Ma and Huang 2023; Wei et al. 2023; Qin et al. 2023; Ryu et al. 2023) have struggled to scale, either due to lack of flexibility or the computational cost of exhaustive search. More recently, diffusion policies (DPs)  (Chi et al. 2023; Wei et al. 2025; He et al. 2025; Wang et al. 2025; Li et al. 2025; Hou et al. 2024; Ze et al. 2024; Wang et al. 2024a; Yang et al. 2024a; Ke, Gkanatsios, and Fragkiadaki 2024; Zhu et al. 2024; Yao et al. 2025; Jia et al. 2025; Cao et al. 2024) have emerged as a promising direction, enabling multi-modal action generation via iterative denoising, and achieving state-of-the-art performance on a range of manipulation benchmarks.
However, existing diffusion policies are typically trained offline and executed without test-time adaptation, using DDPM (Ho, Jain, and Abbeel 2020) or DDIM (Song, Meng, and Ermon 2021) for denoising and treating action generation as an unconstrained stochastic process (see Fig. 1.a). This overlooks key a priori knowledge and task-specific structure available at test time. For instance, the relative pose between the robot and a target object naturally suggests a direction for refinement, while coarse geometric alignment could provide better starting points. Furthermore, standard isotropic Gaussian priors inject noise uniformly across the entire action space, forcing the policy to recover from an uninformative initialization that ignores the specific characteristics of test data, resulting in inefficiencies and increased failure rates.
Figure 1: Comparison of (a) vanilla DPs and (b) our ADP.
@@ 1 Introduction
  • We propose a coarse initialization strategy based on Fast Global Registration (FGR) between test-time point clouds of the gripper and target scene, ensuring the diffusion process begins from a geometrically meaningful configuration.
@@
ADP is training-free and plug-and-play, making it readily deployable with existing pre-trained diffusion models. By implementing ADP with 3D Diffuser Actor (Ke, Gkanatsios, and Fragkiadaki 2024), we create ADPro, which improves sample efficiency, task success, and zero-shot generalization, outperforming state-of-the-art diffusion-based baselines with up to 25% fewer inference steps and 9% points improvement in success rates on long-horizon, contact-rich tasks.
2 Related Work
Diffusion Policies for Robot Control.
@@ 2 Related Work
Diffusion policies (Chi et al. 2023; Wei et al. 2025; He et al. 2025; Wang et al. 2025; Li et al. 2025; Hou et al. 2024; Ze et al. 2024; Wang et al. 2024a; Yang et al. 2024a; Ke, Gkanatsios, and Fragkiadaki 2024; Zhu et al. 2024; Yao et al. 2025; Jia et al. 2025; Cao et al. 2024), a new type of robotic manipulation methods that combine diffusion models (Ho, Jain, and Abbeel 2020; Song, Meng, and Ermon 2021; Song et al. 2021; Li et al. 2023) with imitation learning, have attracted widespread attention for their ability to generate complex and realistic trajectories. These models are trained via behavior cloning to produce action sequences conditioned by visual inputs, robot states, or even language instructions. Existing approaches can be broadly categorized into two main types based on the input data type. The first one relies on 2D observations (Chi et al. 2023; He et al. 2025; Hou et al. 2024; Wang et al. 2024a; Zhu et al. 2024), typically using single or multiple RGB images. Since 2D images are heavily influenced by perspective and texture, such methods tend to be sensitive to viewpoint changes and variations in object color and texture, resulting in limited robustness. The second category leverages 3D visual information (Wei et al. 2025; Wang et al. 2025; Ze et al. 2024; Yang et al. 2024a; Ke, Gkanatsios, and Fragkiadaki 2024; Yao et al. 2025; Jia et al. 2025; Cao et al. 2024), such as point clouds or RGB-D images, offering richer environmental representations and greater generalization to variations in perspective and appearance. However, these methods face two key challenges: (1) Limited generalization to object shapes, positions, and tasks. Most approaches rely solely on gradient estimation from the diffusion model, without incorporating guidance from the target object to inform action updates. (2) Slow inference. The reverse diffusion process typically requires around 100 iterations, limiting the efficiency of action generation. Here, we address both challenges effectively by leveraging guidance from the testing data and applying spherical manifold and initial action constraints.
Robot Policy Generalization.
@@
In robot policy learning, the achievement of generalizable policies has been a long-standing goal (Kaelbling 2020). Two main approaches are commonly used to improve controller generalization: training models on large, diverse datasets, and leveraging the broad generalization capabilities of large vision-language foundation models. For example, some researchers have expanded their training datasets by including more object instances or categories (Finn and Levine 2017; Levine et al. 2018; Depierre, Dellandréa, and Chen 2018; Young et al. 2021; Stone et al. 2023; Wang et al. 2025), novel object combinations (Dasari and Gupta 2021; Jang et al. 2022), diverse language instructions (Jiang et al. 2022; Nair et al. 2022; Wei et al. 2025), and multiple robot embodiments (Vuong et al. 2023). Given the strong representational and generalization capabilities of vision-language models (VLMs), recent works such as RT-2 (Brohan et al. 2023), RoboFlamingo (Li et al. 2024), $\pi$(Black et al. 2024a), ECoT (Zawalski et al. 2024), and Robouniview (Liu et al. 2024) have adopted pre-trained VLMs as the backbone of robotic policies to enhance robustness and generalization. Other research explores alternative perspectives, including domain adaptation (Bousmalis et al. 2018; Fang et al. 2018; Jeong et al. 2020; Ma et al. 2024a; Wu et al. 2025b, a), continuous learning (Yang et al. 2023; Wan et al. 2024; Wang et al. 2024b), and the use of prior constraints (Ma et al. 2024b; Huang et al. 2024; Yao et al. 2025). Despite recent advances, most methods focus on improving policy learning during training, while underutilizing the potential guidance from test-time data during inference. A key challenge lies in how to effectively construct guidance using observational data to improve the generalization of the policy, a problem that remains largely unexplored, and which we aim to address in this work.
3 Method
@@
Action generation inherently involves producing non-Euclidean data, as actions reside within the $\mathrm{SE}(3)$ Lie group. Therefore, integrating a more efficient diffusion policy with a well-designed initial noisy action proposal can greatly enhance generation efficiency. As illustrated in Fig. 2, the proposed adaptive diffusion policy for manipulation action generation accelerates the process through two key improvements. First, an initial action proposal provides the diffusion process with a more reasonable and target-proximal starting action, effectively reducing the number of denoising steps required. Second, manifold constraints guide the denoising process along a more precise trajectory, preventing unnecessary backtracking (see Fig. 4). This results in diffusion trajectories with shorter geodesic distances and fewer diffusion steps, significantly improving overall efficiency.
@@
Figure 2: The pipeline of our adaptive diffusion policy ADPro. We first propose reasonable initial noisy actions $\mathbf{a}^{M}$ with the rough registration module, then, denoise $\mathbf{a}^{M}$ under the guidance of observations and manifold constraint (see Eq. (9) for details).
3.1 Problem Statement.
@@
Given a set of expert demonstrations containing continuous-valued action-observation pairs, denoted as $\mathcal{D}=\{(\mathbf{O}_{k},\mathbf{A}_{k})\}_{k=1}^{N}$, our objective is to fit a visuomotor policy that maps the observation space $\mathbf{O}$ to the action space $\mathbf{A}$. In this work, each observation $\mathbf{O}$ is two posed RGB-D images or point clouds. Each action $\mathbf{A}:=(\mathbf{R},\mathbf{v},w)$ is a 7D pose, defined by position $(x,y,z)$, rotation $(roll,pitch,yaw)$, and the gripper width $w$. Consequently, the pose $(\mathbf{R},\mathbf{v})$ of action lies on an $\mathrm{SE}(3)$ manifold, where $\mathbf{R}\in\mathrm{SO}(3)$ is the rotation matrix, and $\mathbf{v}\in\mathbb{R}^{3}$ is the translation vector.
@@
Let $\mathbf{P}_{k}\in\mathbb{R}^{4\times 4}$ be the current $\mathrm{SE}(3)$ pose of the gripper in the world frame, action $\mathbf{A}_{k}$ specifies a desired pose of the gripper. The pose can be either absolute ($\mathbf{P}_{k+1}=(\mathbf{R}_{k},\mathbf{v}_{k})$, also called position control) or relative ($\mathbf{P}_{k+1}=\mathbf{R}_{k}\mathbf{P}_{k}+\mathbf{v}_{k}$, also called velocity control). To facilitate noise injection and removal using addition and subtraction, as in the standard diffusion process, we vectorize $\mathbf{A}_{k}$ into a vector $\mathbf{a}_{k}$ during diffusion and denoising. After denoising, the noise-free action vectors are converted back into $\mathbf{R}$, $\mathbf{v}$, and $w$ for manipulation.
@@
For trajectory generation, $\mathbf{O}=\{\mathbf{O}_{k-(m-1)},\dots,$ $\mathbf{O}_{k-1},\mathbf{O}_{k}\}$, $\mathbf{a}=\{\mathbf{a}_{k},\mathbf{a}_{k+1},\dots,\mathbf{a}_{k+(n-1)}\}$ where $m$ is the number of history steps observed and $n$ is the number of future action steps. Given that diffusion policies involve both multiple-action trajectories across different positions and a single denoising trajectory for each action at a given position, we introduce a clear notation system to avoid ambiguity. We use a subscript to denote individual action $\mathbf{a}_{k}$ within action trajectories and a superscript to represent noisy action trajectory $\mathbf{a}^{t}$ during denoising process for a single action.
3.2 Preliminaries
Diffusion Policy.
@@ 3.2 Preliminaries
$\mathbf{a}^{t-1}_{k}=\alpha(\mathbf{a}^{t}_{k}-\gamma\varepsilon_{\theta}(\mathbf{O}_{k},\mathbf{a}^{t}_{k},t)+\epsilon),$ (1)
where $\epsilon\sim\mathcal{N}(0,\sigma^{2}I)$. $\alpha,\gamma,\sigma$ are functions of the denoising step $t$ (also known as the noise schedule). The action $\mathbf{a}^{0}$ is expected to be the corresponding action from the expert policy $\pi:\mathbf{O}\mapsto\mathbf{a}$.
Training-Free Guided Diffusion Models.
Classifier guidance (Dhariwal and Nichol 2021) is the first work that utilizes the pre-trained diffusion model for conditional image generation in a training-free fashion. Specifically, considering the Bayes rule $p(x|y)=p(y|x)p(x)/p(y)$, it introduces the given condition with an additional likelihood term $p(x_{t}|y)$:
@@
where $y$ is the class label, $x_{t}\sim\mathcal{N}(\sqrt{\bar{\alpha}_{t}}x_{0},\sqrt{1-\bar{\alpha}_{t}}I)$, $\bar{\alpha}_{t}:=\prod_{s=1}^{t}\alpha_{s}$. Without training the time-dependent classifier to estimate $p(y|x_{t})$, these training-free guided diffusion models (Dhariwal and Nichol 2021; Chung et al. 2023; Yang et al. 2024b) just need a pre-trained diffusion prior $\varepsilon_{\theta}$ and a differentiable loss $L(x_{0},y)$ which is defined on the support of $x_{0}$. They use Tweedie’s formula to calculate $\nabla_{x_{t}}\log p(y|x_{t})$ by estimating $\hat{x}_{0}$ based on $x_{t}$:
$\begin{split}\hat{x}_{0}(x_{t})&\approx\mathbb{E}[x_{0}|x_{t}]=(x_{t}+b_{t}^{2}\nabla_{x_{t}}\log p(x_{t}))/a_{t},\\ \nabla_{x_{t}}\log p(y|x_{t})&\approx\nabla_{x_{t}}\log p(y|\hat{x}_{0}(x_{t}))=\eta\nabla_{x_{t}}L(\hat{x}_{0}(x_{t}),y).\end{split}$ (3)
@@
3.3 Adaptive Diffusion Policy with Manifold Constraint
Guided diffusion models have achieved significant success in generating Euclidean data, such as images. Given the limitations of diffusion policies in generalizing across different tasks, we believe that training-free guided diffusion policies offer an effective solution. However, existing guided diffusion models primarily focus on Euclidean data generation, and in the domain of robotic manipulation, constructing effective guidance remains a challenge.
In robotic manipulation, the vanilla DPs takes $\varepsilon_{\theta}(\mathbf{O}_{k},\mathbf{a}_{k}^{t},t)$ as the denoising direction, serving as the primary update direction for action refinement. However, it does not explicitly leverage the target object to optimize the action more explicitly and effectively. Actually, the relative position between the target object and the end-effector naturally provides the gradient direction for refining the action, directly guiding the end-effector toward objects. Therefore, we integrate the observation guidance $\nabla_{\mathbf{a}}L(\cdot,\cdot)$ into diffusion process and obtain
$\tilde{\mathbf{a}}_{k}^{t-1}$ $=DDPM(\mathbf{a}^{t}_{k},\varepsilon_{\theta}(\mathbf{O}_{k},\mathbf{a}^{t}_{k},t),t)\,$, (5)
$\mathbf{a}_{k}^{t-1}$ $=\tilde{\mathbf{a}}_{k}^{t-1}-\eta\nabla_{\mathbf{a}}L(\tilde{\mathbf{A}}_{k}^{t-1}\mathbf{O}_{k,0},\mathbf{O}_{k,1})\,.$ (6)
Here, $\mathbf{O}_{k,0}$ and $\mathbf{O}_{k,1}$ represent the gripper and scene point clouds extracted from $\mathbf{O}_{k}$, respectively. $\tilde{\mathbf{A}}_{k}^{t-1}$ denotes the matrix form of $\tilde{\mathbf{a}}_{k}^{t-1}$, while $L(\cdot,\cdot)$ is Chamfer distance.
However, since $\tilde{\mathbf{A}}_{k}^{t-1}$ is a noisy action rather than the final action executed by the robot, the guidance provided by Eq. (6) is not sufficiently accurate. Inspired by DDPM (Ho, Jain, and Abbeel 2020), we obtain the noise-free action $\hat{\mathbf{a}}_{k}^{0}$ from $\tilde{\mathbf{a}}_{k}^{t-1}$ using the diffusion function, and reformulate Eq. (6) as
$\hat{\mathbf{a}}_{k}^{0}=(\tilde{\mathbf{a}}_{k}^{t-1}-\sqrt{1-\bar{\alpha}_{t}}\varepsilon_{\theta}(\mathbf{O}_{k},\mathbf{a}^{t}_{k},t))/\sqrt{\bar{\alpha}_{t}}\,$, (7)
$\mathbf{a}_{k}^{t-1}=\tilde{\mathbf{a}}_{k}^{t-1}-\eta\nabla_{\mathbf{a}}L(\tilde{\mathbf{A}}_{k}^{0}\mathbf{O}_{k,0},\mathbf{O}_{k,1})\,.$ (8)

### v2

# ADPro: a Test-time Adaptive Diffusion Policy via Manifold-constrained Denoising and Task-aware Initialization for Robotic Manipulation

@@ Abstract
Index Terms:
diffusion policy, robotics manipulation, test-time adaptation, training-free, task-aware guidance.
I Introduction
Autonomous robotic manipulation in unstructured environments requires policies that not only predict accurate actions but also generalize well across diverse tasks and scenes. Early approaches based on rule-based controllers  [1, 2, 3, 4] or dense discriminative models  [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19] have struggled to scale, either due to lack of flexibility or the computational cost of exhaustive search. More recently, diffusion policies (DPs)  [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34] have emerged as a promising direction, enabling multi-modal action generation via iterative denoising, and achieving state-of-the-art performance on a range of manipulation benchmarks.
However, existing diffusion policies are typically trained offline and executed without test-time adaptation, using DDPM [35] or DDIM [36] for denoising and treating action generation as an unconstrained stochastic process (see Fig. 1.a). This overlooks key a priori knowledge and task-specific structure available at test time. For instance, the relative pose between the robot and a target object naturally suggests a direction for refinement, while coarse geometric alignment could provide better starting points. Furthermore, standard isotropic Gaussian priors inject noise uniformly across the entire action space, forcing the policy to recover from an uninformative initialization that ignores the specific characteristics of test data, resulting in inefficiencies and increased failure rates.
Figure 1: Comparison of (a) vanilla DPs and (b) our ADP. Vanilla DPs generate robot actions by progressively denoising from a random initialization, but this process typically unfolds in an unconstrained Euclidean space. In contrast, our ADP employs task-aware initialization and constrains updates along both task and spherical manifolds, yielding a more generalizable policy.
@@ I Introduction
  • We propose a coarse initialization strategy based on Fast Global Registration between test-time point clouds of gripper and target scene, ensuring the diffusion process begins from a geometrically meaningful configuration.
@@
ADP is training-free and plug-and-play, enabling seamless deployment with existing diffusion policies. Building on the pre-trained 3D Diffuser Actor [29], we developed ADPro, which improves sample efficiency, task success, and zero-shot generalization, outperforming state-of-the-art diffusion-based baselines with up to 25% fewer inference steps and 9% points improvement in success rates on long-horizon tasks.
II Related Work
Diffusion Policies for Robot Control
@@ II Related Work
Diffusion policies [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33], a new type of robotic manipulation methods that combine diffusion models [35, 36, 37, 38] with imitation learning, have attracted widespread attention for their ability to generate complex and realistic trajectories. These models are trained via behavior cloning to produce action sequences conditioned by visual inputs, robot states, or even language instructions. Existing approaches can be broadly categorized into two main types based on the input data type. The first one relies on 2D observations [20, 22, 25, 27, 30], typically using single or multiple RGB images. Since 2D images are heavily influenced by perspective and texture, such methods tend to be sensitive to viewpoint changes and variations in object color and texture, resulting in limited robustness. The second category leverages 3D visual information [21, 23, 26, 28, 29, 31, 32, 33], such as point clouds or RGB-D images, offering richer environmental representations and greater generalization to variations in perspective and appearance. However, these methods face two key challenges: (1) Limited generalization to object shapes, positions, and tasks. Most approaches rely solely on gradient estimation from the diffusion model, without incorporating guidance from the target object to inform action updates. (2) Slow inference. The reverse diffusion process typically requires around 100 iterations, limiting the efficiency of action generation. Here, we address both challenges effectively by leveraging guidance from the testing data and applying spherical manifold and initial action constraints.
Robot Policy Generalization
@@
In robot policy learning, the achievement of generalizable policies has been a long-standing goal [39, 40]. Two main approaches are commonly used to improve controller generalization: training models on large, diverse datasets, and leveraging the broad generalization capabilities of large vision-language foundation models. For example, some researchers have expanded their training datasets by including more object instances or categories [41, 42, 43, 44, 45, 23], novel object combinations [46, 47], diverse language instructions [48, 49, 21, 50], and multiple robot embodiments [51]. Given the strong representational and generalization capabilities of vision-language models (VLMs), recent works such as RT-2 [52], RoboFlamingo [53], $\pi$[54], ECoT [55], and Robouniview [56] have adopted pre-trained VLMs as the backbone of robotic policies to enhance robustness and generalization. Other research explores alternative perspectives, including domain adaptation [57, 58, 59, 60, 61, 62], continuous learning [63, 64, 65], and the use of prior constraints [66, 67, 31]. Despite recent advances, most methods focus on improving policy learning during training, while underutilizing the potential guidance from test-time data during inference. A key challenge lies in how to effectively construct guidance using observational data to improve the generalization, a problem that remains largely unexplored, and which we aim to address in this work.
III Method
@@
Action generation inherently involves producing non-Euclidean data, as actions reside within the $\mathrm{SE}(3)$ Lie group. Therefore, integrating a more efficient diffusion policy with a well-designed initial noisy action proposal can greatly enhance generation efficiency. As illustrated in Fig. 2, the proposed adaptive diffusion policy for manipulation action generation accelerates the process through two key improvements. First, a manifold-constrained denoising – guiding updates along both task and spherical manifolds, effectively enhancing its generalizability and preventing unnecessary backtracking (see Fig. 4). Second, task-aware initialization – driving the diffusion process towards a more accurate direction instead of random noise, effectively reducing the number of denoising steps required. Together, these components transform action generation from blind stochastic sampling into a geometry-aware, adaptive optimization process.
@@
Figure 2: The pipeline of our adaptive diffusion policy ADPro. We first propose reasonable initial noisy actions $\mathbf{a}_{M}$ with the rough registration module, then, denoise $\mathbf{a}_{M}$ under the guidance of observations and manifold constraint (see Eq. (9) for details).
III-A Problem Statement.
@@
Given a set of expert demonstrations containing continuous-valued action-observation pairs, denoted as $\mathcal{D}=\{(\mathbf{O}^{k},\mathbf{A}^{k})\}_{k=1}^{N}$, our objective is to fit a visuomotor policy that maps the observation space $\mathbf{O}$ to the action space $\mathbf{A}$. In this work, each observation $\mathbf{O}$ is two posed RGB-D images or point clouds. Each action $\mathbf{A}:=(\mathbf{v},\mathbf{R},w)$ is a 7D pose, defined by position $(x,y,z)$, rotation $(roll,pitch,yaw)$, and the gripper width $w$. Consequently, the pose $(\mathbf{v},\mathbf{R})$ of action lies on an $\mathrm{SE}(3)$ manifold, where $\mathbf{R}\in\mathrm{SO}(3)$ is the rotation matrix, and $\mathbf{v}\in\mathbb{R}^{3}$ is the translation vector.
@@
Let $\mathbf{P}^{k}\in\mathbb{R}^{4\times 4}$ be the current position of the gripper in the world frame, and action $\mathbf{A}^{k}$ specifies a desired pose of the gripper. The pose can be either absolute ($\mathbf{P}^{k}=(\mathbf{R}^{k},\mathbf{v}^{k})$, also called position control) or relative ($\mathbf{P}^{k}=(\prod_{i=1}^{k}\mathbf{R}^{i}+\sum_{i=1}^{k}\mathbf{v}^{i}$), also called velocity control). To facilitate noise injection and removal using addition and subtraction, as in the standard diffusion process, we vectorize $\mathbf{A}^{k}$ into a vector $\mathbf{a}^{k}$ during diffusion and denoising. After denoising, the noise-free action vectors are converted back into $\mathbf{R}$, $\mathbf{v}$, and $w$ for manipulation.
@@
For trajectory generation, $\mathbf{O}=\{\mathbf{O}^{k-(m-1)},\dots,\mathbf{O}^{k-1},\mathbf{O}^{k}\}$ , $\mathbf{a}=\{\mathbf{a}^{k},\mathbf{a}^{k+1},\dots,\mathbf{a}^{k+(n-1)}\}$ where $m$ is the number of history steps observed and $n$ is the number of future action steps. Since diffusion policies involve both multi-step action trajectories across time and a denoising trajectory for each individual action, we follow standard notation conventions to avoid ambiguity: superscripts denote individual actions $\mathbf{a}^{k}$ within a manipulation trajectory, while subscripts denote the noisy action trajectory $\mathbf{a}_{t}$ during the denoising process of diffusion for a single action.
III-B Preliminaries
Diffusion Policy
@@ III-B Preliminaries
$\mathbf{a}_{t-1}^{k}=\alpha_{t}(\mathbf{a}_{t}^{k}-\gamma_{t}\varepsilon_{\theta}(\mathbf{O}^{k},\mathbf{a}_{t}^{k},t)+\epsilon),$ (1)
where $\epsilon\sim\mathcal{N}(0,\sigma_{t}^{2}I)$. $\alpha_{t},\gamma_{t},\sigma_{t}$ are functions of the denoising step $t$ (also known as the noise schedule). The action $\mathbf{a}_{0}$ is expected to be the corresponding action from the expert policy $\pi:\mathbf{O}\mapsto\mathbf{a}$.
Training-Free Guided Diffusion Models
Classifier guidance [68] is the first work that utilizes the pre-trained diffusion model for conditional image generation in a training-free fashion. Specifically, considering the Bayes rule $p(x|y)=p(y|x)p(x)/p(y)$, it introduces the given condition with an additional likelihood term $p(x_{t}|y)$:
@@
where $y$ is the class label, $x_{t}\sim\mathcal{N}(\sqrt{\bar{\alpha}_{t}}x_{0},(1-\bar{\alpha}_{t})I)$, $\bar{\alpha}_{t}:=\prod_{s=1}^{t}\alpha_{s}$. Without training the time-dependent classifier to estimate $p(y|x_{t})$, these training-free guided diffusion models [68, 69, 70] just need a pre-trained diffusion prior $\varepsilon_{\theta}$ and a differentiable loss $L(x_{0},y)$ which is defined on the support of $x_{0}$. They use Tweedie’s formula to calculate $\nabla_{x_{t}}\log p(y|x_{t})$ by estimating $\hat{x}_{0}$ based on $x_{t}$:
$\begin{split}\hat{x}_{0}(x_{t})\approx\mathbb{E}[x_{0}|x_{t}]&=(x_{t}+\sqrt{1-\bar{\alpha}_{t}}\nabla_{x_{t}}\log p(x_{t}))/\sqrt{\bar{\alpha}_{t}},\\ \nabla_{x_{t}}\log p(y|x_{t})\approx&\nabla_{x_{t}}\log p(y|\hat{x}_{0}(x_{t}))=\eta\nabla_{x_{t}}L(\hat{x}_{0}(x_{t}),y).\end{split}$ (3)
@@
III-C Manifold-Constrained Denoising
Guided diffusion models have shown strong generalization at inference, achieving remarkable success in generating Euclidean data such as images. Yet, most existing approaches remain limited to image domains, and in robotic manipulation, constructing effective guidance remains a key challenge.
In robotic manipulation, vanilla DPs (non-guided) adopt $\varepsilon_{\theta}(\mathbf{O}^{k},\mathbf{a}^{k}_{t},t)$ as the denoising direction, which serves as the primary update for action refinement. Yet, this approach does not explicitly exploit the target object to optimize actions more effectively. To address this, we introduce ADP, which incorporates geometry-aware updates. Specifically, we identify two complementary manifolds that constrain and guide the denoising process: the task manifold and the spherical manifold.
Task manifold
Robot actions lie on the $\mathrm{SE}(3)$ Lie group, where the relative pose between the end-effector and the target object defines a natural gradient direction toward task success. To exploit this structure, we incorporate observation guidance $\nabla_{\mathbf{a}}L(\cdot,\cdot)$ into the diffusion process, steering updates along the geodesic paths of the task manifold. This yields the following guided variant of the diffusion policy:
$\tilde{\mathbf{a}}^{k}_{t-1}$ $=DDPM(\mathbf{a}_{t}^{k},\varepsilon_{\theta}(\mathbf{O}^{k},\mathbf{a}_{t}^{k},t),t)\,$, (5)
$\mathbf{a}^{k}_{t-1}$ $=\tilde{\mathbf{a}}^{k}_{t-1}-\eta\nabla_{\mathbf{a}}L(\tilde{\mathbf{A}}^{k}_{t-1}\mathbf{O}^{k,0},\mathbf{O}^{k,1})\,.$ (6)
Here, $\mathbf{O}^{k,0}$ and $\mathbf{O}^{k,1}$ represent the gripper and scene point clouds extracted from $\mathbf{O}_{k}$, respectively. $\tilde{\mathbf{A}}^{k}_{t-1}$ denotes the matrix form of $\tilde{\mathbf{a}}^{k}_{t-1}$, while $L(\cdot,\cdot)$ is Chamfer distance.
However, since $\tilde{\mathbf{A}}^{k}_{t-1}$ is a noisy action rather than the final action executed by the robot, the guidance provided by Eq. (6) is not sufficiently accurate. Inspired by DDPM [35], we obtain the noise-free action $\hat{\mathbf{a}}^{k}_{0}$ from $\tilde{\mathbf{a}}^{k}_{t-1}$ using the diffusion function, and reformulate Eq. (6) as
$\hat{\mathbf{a}}^{k}_{0}=(\tilde{\mathbf{a}}^{k}_{t-1}-\sqrt{1-\bar{\alpha}_{t}}\varepsilon_{\theta}(\mathbf{O}^{k},\mathbf{a}_{t}^{k},t))/\sqrt{\bar{\alpha}_{t}}\,$, (7)
$\mathbf{a}^{k}_{t-1}=\tilde{\mathbf{a}}^{k}_{t-1}-\eta\nabla_{\mathbf{a}}L(\hat{\mathbf{A}}^{k}_{0}\mathbf{O}^{k,0},\mathbf{O}^{k,1})\,$, (8)
where $\hat{\mathbf{A}}^{k}_{0}$ denotes the matrix form of $\hat{\mathbf{a}}^{k}_{0}$,